Congratulations! You have just won 25 video game tokens to use on these three challenging arcade gamse: Mighty Mediecal Monsters, Radical Racing Robots and Adventures Around the Amazon. All you have to do to collect your prize is answer this question: How many different ways are there to spend your 25 tokens on these three games provided that you play each game at least once?
Answer by khwang(438)

Assume all the 25 tokens are the same, the three games use 25 tokens
without caring the order as repeating combinations 3 different objects
25 times.
Just like the number of positive integer solutions of the equation
X + Y + Z = 25
Or equivalently
X-1 + Y-1 + Z-1 = 22.
Since X-1 >=0, Y-1>=0,and Z-1>=0(means each game at least once)
There are H(3,22) = C(3+22-1,22) = C(24,22) = 24*23/2 = 276...Answer